Decide whether the pair of lines is parallel, perpendicular or neither
3x-4y=-14 and 8x+6y=-13

Answer by stanbon(75887)

Decide whether the pair of lines is parallel, perpendicular or neither
3x-4y=-14 and 8x+6y=-13
--------------
Determine the slope of each linear equation:
3x-4y=-14
4y = 3x+14
y = (3/4)x + (14/4)
slope = 3/4
=====================
8x + 6y = -13
6y = -8x - 13
y = (-8/6)x - (13/6)
slope = -8/6 = -4/3
--------------------------
Since (3/4)(-4/3) = -1
the lines are perpendicular.
==============================
Cheers,
Stan H.